How many grams of aluminum oxide are produced from the synthesis reaction of 8.7 g of Al with oxygen gas? 4Al + 3O2 -> 2Al2O3


1
Expert's answer
2021-01-11T07:17:50-0500

4Al + 3O2 -> 2Al2O3

M (Al) = 27 g/mol

M (Al2O3) = 101.96 g/mol

n (Al) = m/M = 8.7/27 = 0.32 mol

n (Al2O3) = 1/2 x n (Al) = 1/2 x 0.32 = 0.16 mol

m (Al2O3) = 0.16 x 101.96 = 16.4 g